### v1

# Fractal dimensions and profinite groups

Abstract.
Let $T$ be a finitely branching rooted tree such that any node has at least two successors. The path space $[T]$ is an ultrametric space where the distance of distinct paths $f,g$ is $1/|T_{n}|$, with $T_{n}$ denoting the set of nodes at distance $n$ from the root. Let $S$ be a subtree of $T$ without leaves that is level-wise uniformly branching, in the sense that the number of successors of a node only depends on its distance from the root. We show that the Hausdorff and lower box dimensions coincide for $[S]$, and the packing and upper box dimensions also coincide. We give geometric proofs, as well as proofs based on the point to set principles. We use the first result to reprove a result of Barnea and Shalev on the dimension of closed subgroups of a profinite group $G$, referring only on the geometric structure of the closed subgroup in the canonical path space given by an inverse system for $G$.
Mayordomo was supported in part by Spanish Ministry of Science and Innovation grants TIN2016-80347-R and PID2019-104358RB-I00 PID2022-138703OB-I00 and by the Science dept. of Aragon Government: Group Reference T64__\__20R (COSMOS research group).
Nies was supported in part by the Marsden fund of New Zealand, grant UoA-1931
1. Introduction
We study fractal dimensions in the ultrametric setting. Our initial motivation was a description of the Hausdorff dimension of a closed subgroup $U$ of a profinite group $G$ by Barnea and Shalev [2, Thm. 2.4], where the metric is given by an inverse system $(L_{n},p_{n})_{n\in{\mathbb{N}}}$ for $G$. They showed that this dimension equals the lower box dimension:
$\dim_{H}(U)=\liminf_{n}\log|U_{n}|/\log|L_{n}|$,
where $U_{n}$ is the canonical projection of $U$ into $L_{n}$. We will show that this equality can be obtained by an analysis of fractal dimensions for certain closed subsets of path spaces of finitely branching rooted trees, without making reference to the group structure. For the result on closed subsets of path spaces, we give a geometric proof, as well as a proof using the point to set principle in the form given in [6, Thm. 4.1]. In the same two ways, we show a corresponding equality for packing dimension.
1.1. Path spaces
Throughout, $T$ will denote a finitely branching rooted tree such that any node has at least two successors. The set $[T]$ of paths on $T$ is a totally disconnected compact space. The clopen sets are the finite unions of sets $[\sigma]=\{f\colon\sigma\prec f\}$, $\sigma\in T$. The space $[T]$ carries the uniform probability measure $\mu_{T}$ that distributes mass evenly among the paths extending the successors of a node. There is a natural ultrametric on $[T]$: let $d(f,g)$ be the measure of the smallest clopen set containing both $f$ and $g$. Denote by $T_{k}$ the $k$-level of $T$. For distinct $f,g\in[T]$ we have
$d(f,g)=\max\{|T_{n}|^{-1}\colon\,f(n)\neq g(n)\}.$
1.2. Box dimensions of closed subsets of $[T]$
We will study fractal dimensions of closed subsets of $[T]$. The lower and upper box (or Minkowski) dimension of a bounded set $X$ in a metric space $M$ are
(1.1) $\underline{\dim}_{B}(X)$ $=$ $\liminf_{\alpha\to 0^{+}}\frac{\log N_{\alpha}(X)}{\log(1/\alpha)}$
(1.2) $\overline{\dim}_{B}(X)$ $=$ $\limsup_{\alpha\to 0^{+}}\frac{\log N_{\alpha}(X)}{\log(1/\alpha)}$
where $N_{\alpha}(X)$ is the least cardinality of a covering of $X$ with sets of diameter at most $\alpha$. Now let $M=[T]$ for a tree $T$ as above, and write $t_{n}=|T_{n}|$. A closed subset of $[T]$ has the form $[S]$ for a unique subtree $S$ of $T$. If $t_{n-1}^{-1}>\alpha\geq t_{n}^{-1}$, then we may assume that the sets in a covering as above are of the form $[\sigma]$ for $|\sigma|=n$, because each set of diameter at most $\alpha$ is contained in such a set $[\sigma]$. So, we have $N_{\alpha}(X)=|S_{n}|$. Since $-\log\alpha\leq\log t_{n}$,
(1.3) $\underline{\dim}_{B}([S])$ $=$ $\liminf_{n\to\infty}\frac{\log|S_{n}|}{\log|T_{n}|}$
(1.4) $\overline{\dim}_{B}([S])$ $=$ $\limsup_{n\to\infty}\frac{\log|S_{n}|}{\log|T_{n}|}.$
1.3. Box dimensions versus Hausdorff/packing dimension
By $V^{*}$ one denotes the set of tuples over a set $V$, seen as a rooted tree where the successor relation is given by appending one element. A subtree of $V^{*}$ is a nonempty subset closed under initial segments. We will always view $T$ as a subtree of ${\mathbb{N}}^{*}$. Thus, $T_{k}$ consists of the strings in $T$ that have length $k$.
Definition 1.1.
We say that a tree $S\subseteq{\mathbb{N}}^{*}$ is level-wise uniformly branching if for each $n$, each $\sigma\in S_{n}$ has the same number of successors on $S$. Note that for such a tree, $\mu_{S}([\sigma])=|S_{n}|^{-1}$ for each $\sigma\in S_{n}$.
Example 1.2.
Let $T=\{0,1\}^{*}$. Let $S$ be a level-wise uniformly branching subtree of $T$ such that for each even $k$, for each $n\in[4^{k},4^{k+1})$, strings at level $n$ have one successor, and for each odd $k$, for each $n\in[4^{k},4^{k+1})$, strings at level $n$ have two successors. Then $\underline{\dim}_{B}([S])\leq 1/4$ and $\overline{\dim}_{B}([S])\geq 3/4$.
For background on Hausdorff dimension and also packing dimension $\dim_{P}$, see for instance the first two chapters of [3]. Here we merely review the pertinent definitions. Let $X$ be a bounded set in a metric space $M$. For $r\geq 0$, by $H^{r}(X)$ one denotes the $r$-dimensional Hausdorff measure
(1.5) $H^{r}(X)=\lim_{\alpha\to 0^{+}}(\inf\sum_{i\in I}\operatorname{diam}(E_{i})^{r}),$
where the infimum is taken over all covers $(E_{i})_{i\in I}$ of $X$ by sets of diameter at most $\alpha$. By definition, $\dim_{H}X$ is the supremum of the $r$ such that $H^{r}(X)>0$. The packing dimension $\dim_{P}([X])$ is defined “from the inside” via disjoint sets of balls with centre in $X$; we refer to [3] for the formal definition. It is known that $\dim_{H}(X)\leq\underline{\dim}_{B}(X)$ and $\dim_{P}(X)\leq\overline{\dim}_{B}(X)$ for any bounded subset $X$ of a metric space $M$ [3, p. 72].
2. Coincidences of dimensions
Let $S$ be a level-wise uniformly branching subtree of $T$. View $[S]$ as a subset of the metric space $[T]$. We show that the Hausdorff and lower box dimensions of $[S]$ coincide, and the packing and upper box dimensions also coincide. This section gives direct proofs of these facts. The final section provides alternative proofs using an algorithmic theory, in particular Martin-Löf randomness on $[S]$ and the point-to-set principles in metric spaces of Lutz et al. [6]. These proofs directly obtain the equalities from the point-to-set principles, rather than relying on the known inequalities mentioned at the end of the previous section.
Theorem 2.1.
Let $S\subseteq T\subseteq{\mathbb{N}}^{*}$ be trees such that each node on $T$ has at least two successors. Suppose that $S$ is level-wise uniformly branching. Then in the metric space $[T]$ we have
$\dim_{H}[S]$ $=$ $\underline{\dim}_{B}[S],\text{ and}$
$\dim_{P}[S]$ $=$ $\overline{\dim}_{B}[S].$
Proof.
It suffices to show the inequalities $\geq$”, as the converse inequalities always hold.
1. For the inequality $\dim_{H}[S]\geq\underline{\dim}_{B}[S]$, we may assume $\underline{\dim}_{B}[S]>0$. It suffices to show that $H^{r}[S]\geq 1$ for each $r<\underline{\dim}_{B}[S]$ such that $r>0$.
We first provide a reduction of the type of coverings we need to consider. Write $t_{n}=|T_{n}|$ and $s_{n}=|S_{n}|$. Suppose we have a covering $(E_{i})_{i\in I}$ of $[S]$ by sets of diameter at most $\alpha$, where $t_{n-1}^{-1}>\alpha\geq t_{n}^{-1}$. In evaluating (1.5) we can assume that $\alpha=t_{n}^{-1}$. The $r$-cost of a family $(E_{i})_{i\in I}$ is by definition
$V_{r}(E_{i})_{i\in I}=\sum_{i\in I}\operatorname{diam}(E_{i})^{r}$.
In evaluating (1.5) we can further assume that each $E_{i}$ intersects $[S]$. Also, without incurring a larger cost we can replace a covering by a covering with sets that contain the original sets but have no larger diameter. So we can assume that each $E_{i}$ is of the form $[\sigma]$ for some $\sigma\in S$, where $|\sigma|\geq n$. Furthermore, since $[S]$ is compact and each set of the form $[\sigma]$ is open, we can now assume that $I$ is finite. To summarise, in evaluating (1.5), the only covers of $[S]$ we need to consider are are covers $\{[\sigma]\colon\sigma\in F\}$ where $F\subseteq S$ is a finite prefix-free set of strings of length at least $n$. We will identify such a set $F\subseteq S$ with the covering it describes, and in particular write $V_{r}(F)$ for the $r$-cost of the covering. So $V_{r}(F)=\sum_{\rho\in F}t_{|\rho|}^{-r}$.
Claim 2.2.
Given $n$, let $F\subseteq S$ be as above. Then $V_{r}(F)\geq s_{k}t_{k}^{-r}$ for some $k\geq n$.
To verify this, we keep replacing $F$ by a covering $F^{\prime}$ of $[S]$ of no larger $r$-cost, until we reach a covering of the form $\{[\sigma]\colon\sigma\in S_{k}\}$ for some $k\geq n$, which then has the required $r$-cost $s_{k}t_{k}^{-r}$.
Let $k\geq n$ be the least length of a string in $F$. For $\sigma\in S_{k}$, write
$v(\sigma)=V_{r}\{\rho\in F\colon\,\rho\succeq\sigma\}.$
Note that $V_{r}(F)=\sum_{\sigma\in S_{k}}v(\sigma)$. Pick $\sigma\in S_{k}$ such that $v(\sigma)$ is least.
  • (a)
    If $\sigma\in F$, then $V_{r}(\{\sigma\})=t_{k}^{-r}$, so $V_{r}(F)\geq s_{k}t_{k}^{-r}$, so we end the process.
  • (b)
    Otherwise, let $F^{\prime}=\{\eta\tau\colon\,\eta\in S_{k}\,\land\,\sigma\tau\in F\}$. Since $S$ is level-wise uniformly branching, we have $F^{\prime}\subseteq S$. Also, $V_{r}(F^{\prime})\leq V_{r}(F)$, each string in $F^{\prime}$ has a length greater than $k$, $F^{\prime}$ is prefix-free, and the maximum length of strings in $F^{\prime}$ is at most the one of $F$. So we can continue the process with $F^{\prime}$ instead of $F$.
After finitely many steps the process ends at (a), delivering the cover of $[S]$ of the required form. This shows the claim.
Since $r<\underline{\dim}_{B}[S]$, there is $K\in{\mathbb{N}}$ such that $\log s_{k}/\log t_{k}\geq r$ for each $k\geq K$, and hence $s_{k}t_{k}^{-r}\geq 1$. For $\alpha<t_{K}^{-1}$, if $(E_{i})_{i\in I}$ is a cover of $[S]$ such that $\operatorname{diam}(E_{i})\leq\alpha$ for each $i\in I$, then by the reductions and the claim we have $V_{r}(E_{i})_{i\in I}\geq s_{k}t_{k}^{-r}$ for some $k\geq K$, and hence $V_{r}(E_{i})_{i\in I}\geq 1$. Thus $H^{r}[S]\geq 1$ as required.
2. $\dim_{P}[S]\geq\overline{\dim}_{B}[S]$. Our proof is based on a result of Tricot [11] (also see [3, Lemma 2.8.1(i)]). By this result,
$\dim_{P}[S]\geq\inf\{\overline{\dim}_{B}(V)\colon V\neq\emptyset\text{ open in}[S]\};$
By the monotonicity of $\overline{\dim}_{B}$, without changing the value we can take the infimum over all sets $V$ of the form $[\sigma]\cap[S]$ where $\sigma\in S$. Fix such a $\sigma$, and write $r=|\sigma|$. Since $S$ is level-wise uniformly branching, for $n\geq r$ the $n$-th level of the tree for $[\sigma]\cap[S]$ has $s_{n}/s_{r}$ elements. So, using (1.4) we have $\overline{\dim}_{B}(V)=\limsup_{n}\frac{\log(s_{n}/s_{r})}{\log t_{n}}=\limsup_{n}\frac{\log(s_{n})}{\log t_{n}}=\overline{\dim}_{B}[S]$. ∎
The following example shows that some hypothesis on $S$, such as being level-wise uniformly branching, is necessary in both equalities of Theorem 2.1.
Example 2.3.
Let $T=2^{<\omega}$ be the full binary tree. For a string $\sigma\in T$, let $k_{\sigma}=\min\{i\colon\,i\leq|\sigma|\lor\sigma_{i}=1\}$. Let
$S=\{\sigma\in T\colon\,\sigma_{r}=0\text{ for each }r\geq 2k_{\sigma}\}$.
Clearly $[S]$ is countable, and hence $\dim_{H}[S]=\dim_{P}[S]=0$. In contrast, since $2\cdot 2^{n/2}\geq|S_{n}|\geq 2^{n/2}$, we have $\underline{\dim}_{B}[S]=\overline{\dim}_{B}[S]=1/2$.
3. Dimensions of closed subgroups of profinite groups
Let the topological group
[…]
be the inverse limit of an inverse system of discrete finite groups, where the maps $p_{n}\colon L_{n+1}\to L_{n}$ are epimorphisms. The group $G$ is compact and has a compatible ultrametric, where the distance between distinct elements $x,y$ of $G$ is $1/|L_{n}|$ for the largest $n$ such that the natural projections of $x$ and $y$ into $L_{n}$ are equal. Barnea and Shalev [2, Thm. 2.4] show that the Hausdorff dimension of a closed subgroup $U$ of $G$ equals its lower box dimension. They rely on Abercrombie [1, Prop. 2.6] for the nontrivial part that the box dimension is no larger than the Hausdorff dimension. Abercrombie proved this via means specific to profinite groups, such as generation of subgroups and the Haar measure. We will show in Theorem 3.1 that this is a purely geometric fact that does not rely on the group structure, but is merely exploiting via Theorem 2.1 the regularity of a description of $U$ as a subtree of the tree given by the inverse system. We also obtain the dual fact that the packing dimension of a closed subgroup equals its upper box dimension.
We may assume $L_{0}$ is the trivial group. Define a tree $T$, where the $n$-th level is $T_{n}=L_{n}$. A string $\eta\in T_{n+1}$ is a successor of $\sigma\in T_{n}$ if $p_{n}(\eta)=\sigma$. Each $\sigma\in T_{n}$ has $|L_{n+1}|/|L_{n}|$ successors. So $T$ is level-wise uniformly branching. With some encoding of the elements of each $L_{n}$ by the first $|L_{n}|$ natural numbers, we can view $T$ as a subtree of ${\mathbb{N}}^{*}$. There is a canonical isometry between $G$ and $[T]$ sending an element $x$ of $G$ to the path $f_{x}$ with $f_{x}(n)=r$ if the natural projection map $G\to L_{n+1}$ sends $x$ to $r$. (To see that the subscript $n+1$ in $L_{n+1}$ is correct, keep in mind that $L_{0}$ is the trivial group. )
[…]
The following is now immediate from Theorem 2.1.
Theorem 3.1 ([2], Thm. 2.4 for the first line of equations).
In the notation above,
$\dim_{H}(U)=\underline{\dim}_{B}(U)=\liminf_{n}\frac{\log|U_{n}|}{\log|L_{n}|}.$
$\dim_{P}(U)=\overline{\dim}_{B}(U)=\limsup_{n}\frac{\log|U_{n}|}{\log|L_{n}|}.$
We end the section with some discussion of fractal dimensions, in particular in profinite groups. The following example shows that $\underline{d}im_{B}(U)<\overline{\dim}_{B}(U)$ is possible.
Example 3.2 (essentially [2], Lemma 4.1).
Let $E$ be the Cantor space $C_{2}^{\omega}$ with symmetric difference $\Delta$ as the group operation. Let the inverse system be given by $L_{n}=C_{2}^{n}$, with the canonical map $p_{n}$ that omits the first bit in a binary string of length $n+1$. For each pair of reals $\alpha,\beta$ with $0\leq\alpha\leq\beta\leq 1$ there is a closed subgroup $U$ of $G$ such that $\underline{\dim}_{B}(U)=\alpha$ and $\overline{\dim}_{B}(U)=\beta$.
Proof.
Let $R\subseteq{\mathbb{N}}$ be a set with lower [upper] density $\alpha$ [$\beta$]. Let $H$ be the subgroup of sequences that have a 00 in positions outside $R$. We have $|S_{n}|=2^{|X\cap n|}$, $|T_{n}|=2^{n}$. Thus $\underline{\dim}_{B}(U)=\alpha$ and $\overline{\dim}_{B}(U)=\beta$ by (1.1) and (1.2), respectively. ∎
If $G$ is a $p$-adic analytic pro-$p$ group [4], then by [2, Thm 1.1],
$\dim_{H}(U)=\dim_{\text{manifold}}(U)/\dim_{\text{manifold}}(G)$,
where $\dim_{\text{manifold}}(X)$ is the dimension of a closed set $X$ in the $p$-adic manifold underlying $G$. Barnea and Shalev [2, end of Section 3] assert that the proof also works for the upper box dimension in place of $\dim_{H}$, so the same expression holds. Thus all the fractal dimensions agree for closed subgroups of $G$; in particular, $\dim_{H}$ and $\dim_{P}$ agree.
The spectrum of a profinite group is the set of Hausdorff dimensions of closed subgroups. The group $E$ given in Example 3.2 has the largest possible spectrum $[0,1]$. Note that $E$ is not (topologically) finitely generated. Whether the spectrum of a finitely generated pro-$p$ group can be $[0,1]$ remains open; see [2, Problem 5]. For an example that comes somewhat close, consider the pro-$p$ group $\mathrm{SL}_{2}^{1}(F_{p}[[t]])$ consisting of matrices such that off diagonal elements are in the maximal ideal $tF_{p}[[t]]$; this is the first congruence subgroup of $\mathrm{SL}_{2}(F_{p}[[t]])$. In case that $p\neq 2$, it has spectrum $[0,2/3]\cup\{1\}$ by [2, Cor 1.5] and the discussion thereafter. It is finitely generated by [10, Prop. 4.11].
4. Algorithmic dimension of points
The remainder of the paper is devoted to the algorithmic theory of fractal dimensions. This present section reviews the Kolmogorov complexity of points in metric spaces, and describes the point-to-set principle for both Hausdorff and packing dimension in that context. We mostly follow Lutz, Lutz and Mayordomo [6]. However, we restrict ourselves to the standard gauge family $\theta_{s}(\alpha)=\alpha^{s}$, which has a precision family in their sense [6, Observation 2.1]. For background also see [8].
Convention 4.1.
All logarithms will be taken with respect to base $2$.
Suppose we are given a separable metric space $M$, together with a dense set $D\subseteq M$ and an onto function $r\colon\{0,1\}^{*}\to D$. For $w\in D$ one lets
(4.1) $C^{M,r}(w)=\min\{C(u)\colon r(u)=w\},$
where $C(u)$ is the usual Kolmogorov complexity of a binary string $u$. (Li and Vitanyi [5] is a standard reference for Kolmogorov complexity.)
Definition 4.2 ([6], (3.1) and (3.2)).
For $x\in M$ and any $\alpha>0$, one defines the Kolmogorov complexity of a point $x$ at precision $\alpha$ by
(4.2) $C_{\alpha}(x)=\min\{C^{M,r}(w)\colon w\in D\,\land\,d(x,w)\leq\alpha\}.$
The algorithmic dimension of $x\in M$ is
(4.3) $\textit{dim}(x):=\liminf_{\alpha\to 0^{+}}C_{\alpha}(x)/\log(1/\alpha).$
The strong algorithmic dimension of $x\in M$ is
(4.4) $\textit{Dim}(x):=\limsup_{\alpha\to 0^{+}}C_{\alpha}(x)/\log(1/\alpha).$
These definitions can be readily relativised to an oracle set $A\subseteq{\mathbb{N}}$: in (4.1), replace $C(u)$ by $C^{A}(u)$, the Kolmogorov complexity of a string $u$ as measured by a universal machine equipped with $A$ as an oracle. Two point-to-set principles [6, Thms 4.1,4.2] determine the Hausdorff and packing dimension of a set $X\subseteq M$ in terms of the relativised algorithmic dimensions of the points in it:
(4.5) $\dim_{H}(X)$ $=$ $\min_{A}\sup_{x\in X}\textit{dim}^{A}(x),$
(4.6) $\dim_{P}(X)$ $=$ $\min_{A}\sup_{x\in X}\textit{Dim}^{A}(x)$
Definition 4.3.
Fix some computable injection $\pi\colon{\mathbb{N}}^{*}\to\{0,1\}^{*}$, say
$(n_{0},\ldots,n_{k-1})\mapsto 0^{n_{0}}1\ldots 0^{n_{k-1}}1$.
The plain descriptive (Kolmogorov) complexity $C(y)$ of $y\in{\mathbb{N}}^{*}$ is defined to be $C(\pi(y))$; in a similar way one defines $K(y)$, $C(y\mid n)$, etc.
The choice of $\pi$ will only affect these values up to a constant. It will be clear from the context whether $C(y)$ is taken to be evaluated for $y\in{\mathbb{N}}^{*}$ or for $y\in\{0,1\}^{*}$. There are now two interpretations of the expression $C(y)$ for a binary string $y$; however, they agree up to a constant.
Recall that throughout, $T\subseteq{\mathbb{N}}^{*}$ is a finitely branching tree such that each node has at least two successors. By $T_{n}$ we denote the set of strings in $T$ of length $n$. The space $[T]$ carries the natural ultrametric
$d(f,g)=\max\{|T_{n}|^{-1}\colon\,f(n)\neq g(n)\}$
where $\max\emptyset=0$ by convention. For real-valued expressions $E,F$ depending on the same parameters, we write $E\leq^{+}F$ for $E\leq F+\mathcal{O}(1)$, and $E=^{+}F$ for $E\leq^{+}F\leq^{+}E$.
Suppose $[T]$ is computable. We describe a computable structure on the metric space $[T]$. Let $\underline{r}\colon\{0,1\}^{*}\to T$ be the computable bijection that is an isomorphism between the length lexicographical orderings on $\{0,1\}^{*}$ and $T$. Given a binary string $u$, let $r(u)$ be a path on $[T]$ extending the string $\underline{r}(u)$. Since $T$ is computable, one can choose $r(u)$ to be uniformly computable in $u$.
We next relate $C_{\alpha}(x)$, the Kolmogorov complexity of a point at precision $\alpha$ in (4.2), to the initial segment complexity of paths on $[T]$.
Proposition 4.4.
Suppose that $T\subseteq{\mathbb{N}}^{*}$ is a computable tree such that each node has at least two successors. Let $r\colon\{0,1\}^{*}\to[T]$ define a computable structure as above. Let $f\in[T]$. For each $\alpha$ such that $|T_{n-1}|^{-1}>\alpha\geq|T_{n}|^{-1}$, we have
$C(f\!\upharpoonright_{n})-2\log n\leq^{+}C(f\!\upharpoonright_{n}\mid n)\leq^{+}C_{\alpha}(f)\leq^{+}C(f\!\upharpoonright_{n})$.
Proof.
The leftmost inequality is standard [5]. For the middle inequality, suppose $w\in D$ and $d(w,f)\leq\alpha$. Given $u\in\{0,1\}^{*}$ such that $r(u)=w$ and $n\in{\mathbb{N}}$, we can compute $w\!\upharpoonright_{n}$, which has to equal $f\!\upharpoonright_{n}$ since $|T_{n-1}|^{-1}>\alpha$. Thus $C(f\!\upharpoonright_{n}\mid n)\leq^{+}C(w)$.
For the rightmost inequality, given a string $y=f\!\upharpoonright_{n}\in T$ we compute the string $u\in\{0,1\}^{*}$ such that $\underline{r}(u)=y$. Since $d(f,r(u))\leq\alpha$, this shows $C_{\alpha}(f)\leq^{+}C(u)=^{+}C(y)$. ∎
The algorithmic dimensions $\textit{dim}(f)$ and $\textit{Dim}(f)$ (see Def. 4.2) of a path $f\in[T]$ can be expressed in terms of the descriptive complexity of its initial segments. This is well-known at least for the case of a binary tree $T$, going back to [7].
Proposition 4.5.
Let $T$ be a computable tree such that each string on $T$ has at least two successors on $T$. For any $f\in[T]$ we have
$\textit{dim}(f)=\liminf_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n}|}=\liminf_{n}\frac{K(f\!\upharpoonright_{n})}{\log|T_{n}|}$
$\textit{Dim}(f)=\limsup_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n}|}=\limsup_{n}\frac{K(f\!\upharpoonright_{n})}{\log|T_{n}|}$
Proof.
We only verify the first assertion; the argument for the second assertion is almost identical. By hypothesis we have $\log|T_{n}|\geq n$, and so $\lim_{n}\log n/\log|T_{n}|=0$. For the left equality, using Proposition 4.4 we have
$\dim(f)=\liminf_{\alpha\to 0^{+}}C_{\alpha}(f)/\log(1/\alpha)=\liminf_{n\to\infty}C(f\!\upharpoonright_{n})/\log|T_{n}|$.
For the right equality, note that for each $n$ and each $y\in T_{n}$, we have $C(y)\leq^{+}K(y)\leq^{+}C(y)+K(n)\leq^{+}C(y)+2\log n$. ∎
5. Martin-Löf randomness and Levin-Schnorr Theorem on $[S]$.
As mentioned, for a computable finitely branching tree $S$ without leaves, the path space $[S]$ carries the uniform measure $\mu_{S}$ determined by the conditions that $\mu_{S}[S]=1$ and $\mu_{S}([\sigma\,\widehat{\ }\,i]\cap[S])=\mu_{S}([\sigma]\cap[S])/k_{\sigma}$ where $k_{\sigma}$ is the number of successors of $\sigma$ on $S$. If $S$ and the function $\sigma\mapsto k_{\sigma}$ are computable, then $\mu_{S}$ is computable. Thus, the usual algorithmic test notions such as Martin-Löf tests (ML-tests) can be used for $[S]$ in place of Cantor space. Also, one can generalise the Levin-Schnorr Theorem (see, e.g., [9, 3.2.9]) to $[S]$. For $X\subseteq S$ let $[X]^{\prec}_{S}=\{f\in S\colon\exists n\,f\!\upharpoonright_{n}\in X\}$ be the open subset of $[S]$ described by $X$. Let
$R_{S,b}=[\{\sigma\in S\colon K(\sigma)\leq-\log\mu_{S}[\sigma]-b\}]^{\prec}_{S}.$
Our version of the Levin-Schnorr theorem is as follows.
Lemma 5.1.
$(R_{S,b})_{b\in{\mathbb{N}}}$ is a universal Martin-Löf test on $[S]$. Thus, it is a ML-test, and $f\in[S]$ is ML-random iff there is $b\in{\mathbb{N}}$ such that
$K(f\!\upharpoonright_{n})>-\log\mu_{S}[f\!\upharpoonright_{n}]-b$ for each $n\in{\mathbb{N}}$.
Proof.
Clearly the sets $R_{S,b}$ are $\Sigma^{0}_{1}$ uniformly in $b$. Given $b$, let $D$ be the set of prefix minimal strings $\sigma\in S$ such that $K(\sigma)\leq-\log\mu([\sigma])-b$. Then
$1\geq\sum_{\sigma\in D}2^{-K(\sigma)}\geq\sum_{\sigma\in D}2^{\log\mu([\sigma])+b}.$
This implies that $\mu_{S}[D]^{\prec}_{S}\leq 2^{-b}$.
Universality is shown as in the proof for Cantor space $2^{{\mathbb{N}}}$; see [9, 3.2.9]. To adapt that proof, one first needs to write down a machine existence theorem for machines with binary input that output strings in ${\mathbb{N}}^{*}$. As this is well-known in principle, we omit the details. ∎
While we are mainly interested in the case that $[S]$ is uncountable, the argument above also works for countable $[S]$. In that case only the empty set is null, so every path is ML-random on $[S]$.
6. Algorithmic dimension versus box dimension
We show that in the case that a computable subtree $S$ of $T$ is level-wise uniformly branching, the lower/upper box dimension of $[S]$ equals the (strong) algorithmic dimension of a ML-random $f\in[S]$. To provide the desired proof of Theorem 2.1 using the point-to-set principles, this fact will be used in relativised form.
Proposition 6.1.
(i) Let $T$ be a computable tree as above, and let $S$ be a computable subtree of $T$. For each $f\in[S]$, we have
$\textit{dim}(f)$$\leq$ $\underline{\dim}_{B}([S]),\text{ and}$
$\textit{Dim}(f)$$\leq$ $\overline{\dim}_{B}([S]).$
(ii) Suppose $S$ is level-wise uniformly branching. Then equalities hold in the case that $f$ is Martin-Löf random in $[S]$ with respect to $\mu_{S}$.
Proof.
We use Proposition 4.5 for both parts. Let $S_{\leq n}=\bigcup_{k\leq n}S_{k}$.
(i) Let $\eta_{S}\colon S\to\{0,1\}^{*}$ be the (computable) isomorphism between the length-lexicographical orderings on $S$ and on $\{0,1\}^{*}$. Recall that to define $C(y)$ and $K(y)$ for $y\in{\mathbb{N}}^{*}$, we fixed a computable injection $\pi\colon{\mathbb{N}}^{*}\to\{0,1\}^{*}$. Also recall that logarithms are taken in base $2$.
Claim. Let $y\in S_{n}$. Then $C(y)\leq^{+}\log|S_{\leq n}|$.
To verify this, let $x=\eta_{S}(y)$. We have $|x|\leq\lceil\log|S_{\leq n}|\rceil$. Since $x$ is binary, $C(x)\leq^{+}|x|$. Since the function $\pi\circ\eta_{S}^{-1}$ is computable, we have $C(y)=C(\pi(\eta_{S}^{-1}(x)))\leq^{+}C(x)$. The claim follows.
To verify the inequalities (i), let $y=f\!\upharpoonright_{n}$. Note that
$\log|S_{\leq n}|\leq\log(n\cdot|S_{n}|)\leq^{+}\log n+\log|S_{n}|$.
So by the claim, $C(y)\leq^{+}\log n+\log|S_{n}|$. Using that $\log n=o(\log|T_{n}|)$,
$\dim(f)$ $=$ $\liminf_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n}|}$
$\leq$ $\liminf_{n}\frac{\log n+\log|S_{n}|}{\log|T_{n}|}=\dim_{B}[S].$
(ii) If $f$ is ML random then $f$ passes the universal test $(R_{S,b})_{b\in{\mathbb{N}}}$ in Lemma 5.1. Thus, using that $S$ is level-wise uniformly branching, we have
$\log|S_{n}|=^{+}-\log\mu_{S}[f\!\upharpoonright_{n}]\leq^{+}K(f\!\upharpoonright_{n})$.
The equalities now follow from Proposition 4.5. ∎
Proof of Th. 2.1 using the point-to-set principle.
Suppose that $E\subseteq{\mathbb{N}}$ is an oracle such that $S\leq_{T}E$, $T\leq_{T}E$, and $E$ can compute from $n$ the number of successors for $S$ at level $n$. Note that the uniform measure $\mu_{S}$ on $[S]$ is $E$-computable.
We show $\dim_{H}[S]=\underline{\dim}_{B}[S]$. By the point-to-set principle for Hausdorff dimension (4.5), it suffices to show that $\min_{A}\sup_{f\in[S]}\textit{dim}^{A}(f)=\underline{\dim}_{B}([S])$. We have the inequality $\leq$” by Prop. 6.1(i) relativised to $E$. For the converse inequality, note that in taking the minimum on the left hand side, we may assume that the oracle $A$ computes $E$, because replacing $A$ by $A\oplus E$ does not increase $\textit{dim}^{A}(f)$. For each such $A$, if we choose $f$ to be MLA-random in $[S]$, by Prop. 6.1(ii) we have $\textit{dim}^{A}(f)=\underline{\dim}_{B}(S)$.
The second equality in Th. 2.1 is obtained in a similar fashion, using the point-to-set principle for packing dimension (4.6) instead. ∎