### v3

# Fractal dimensions and profinite groups

Abstract.
Let $T$ be a finitely branching rooted tree such that any node has at least two successors. The path space $[T]$ is an ultrametric space: for distinct paths $f,g$ let $d(f,g)=1/|T_{n}|$, where $T_{n}$ denotes the $n$-th level of the tree, and $n$ is largest such that $f(n)=g(n)$. Let $S$ be a subtree of $T$ without leaves that is level-wise uniformly branching, in the sense that the number of successors of a node only depends on its level. We show that the Hausdorff and lower box dimensions coincide for $[S]$, and the packing and upper box dimensions also coincide. We give geometric proofs, as well as proofs based on the point-to-set principles. We use the first result to reprove a theorem of Barnea and Shalev on the Hausdorff dimension of closed subgroups of a profinite group $G$, referring only on the geometric structure of the closed subgroup in the canonical path space given by an inverse system for $G$. We obtain an analogous theorem for the packing dimension.
Mayordomo was supported in part by Spanish Ministry of Science and Innovation grants TIN2016-80347-R and PID2019-104358RB-I00 PID2022-138703OB-I00 and by the Science dept. of Aragon Government: Group Reference T64_\_20R (COSMOS research group).
Nies was supported in part by the Marsden fund of New Zealand, grant UoA-1931
1. Introduction
We study fractal dimensions in the ultrametric setting. Our initial motivation was a description of the Hausdorff dimension of a closed subgroup $U$ of a profinite group $G$ by Barnea and Shalev [2, Thm. 2.4], where the metric is given by an inverse system $(L_{n},p_{n})_{n\in{\mathbb{N}}}$ for $G$. They showed that this dimension equals the lower box dimension:
$\dim_{H}(U)=\liminf_{n}\log|U_{n}|/\log|L_{n}|$,
where $U_{n}$ is the canonical projection of $U$ into $L_{n}$. We will show that this equality can be obtained without making reference to the group structure. Instead, one can analyse the fractal dimensions for closed subsets of path spaces of finitely branching rooted trees that are given by level-wise uniformly branching subtrees.
For the result on closed subsets of path spaces, we give a geometric proof, as well as a proof using the point-to-set principle in the form given in [9, Thm. 4.1], which determines the dimension of a set via an algorithmic notion of dimension for individual points. In the same two-pronged manner we analyse the packing dimension of such sets, and apply the result to profinite groups as well. We include the second approach for three reasons. Unlike the geometric proofs, the proofs based on point-to-set principles use no preliminary results from the theory of fractional dimensions: neither the “easy” inequalities (1.8), nor Tricot’s result [14] on packing dimension. Also, this line of argument promises to be extendable to settings less restrictive than the level-wise uniformly branching tree. Finally, this approach leads to a result of interest on its own (Prop. 6.1) on Martin-Löf random paths on computable trees.
1.1. Path spaces
Throughout, $T$ will denote a finitely branching rooted tree such that any node has at least two successors. For $n\geq 0$ we denote the $n$-level of $T$ by $T_{n}$. By a path of $T$ one means a sequence $f=(a_{0},a_{1},\ldots)$ such that $a_{n}\in T_{n}$ and $a_{n+1}$ is a successor of $a_{n}$, for each $n$. The set $[T]$ of paths on $T$ is a totally disconnected compact space. The clopen sets are the finite unions of sets $\sigma]=\{f\colon f(n)=\sigma\}$, where $\sigma\in T_{n}$.
There is a natural ultrametric $d$ on $[T]$: For distinct $f,g\in[T]$,
(1.1) $d(f,g)=\min\{1/|T_{n}|\colon\,f(n)=g(n)\}.$
Since every node $\sigma$ on $T$ has at least two successors, the set $\sigma]$ has diameter $1/|T_{n}|$ for $\sigma\in T_{n}$.
1.2. Box dimensions of closed subsets of $[T]$
We will study fractal dimensions of closed subsets of $[T]$. The lower and upper box (or Minkowski) dimension of a bounded set $X$ in a metric space $M$ are
(1.2) $\underline{\dim}_{B}(X)$ $=$ $\liminf_{\alpha\to 0^{+}}\frac{\log N_{\alpha}(X)}{\log(1/\alpha)}$
(1.3) $\overline{\dim}_{B}(X)$ $=$ $\limsup_{\alpha\to 0^{+}}\frac{\log N_{\alpha}(X)}{\log(1/\alpha)}$
where $N_{\alpha}(X)$ is the least cardinality of a covering of $X$ with sets of diameter at most $\alpha$. Now let $M=[T]$ for a tree $T$ as above, and write $t_{n}=|T_{n}|$. A closed subset $X$ of $[T]$ has the form $[S]$ for a unique subtree $S$ of $T$ without leaves. If $t_{n-1}^{-1}>\alpha\geq t_{n}^{-1}$, then we may assume that the sets in a covering as above are of the form $\sigma]$ for $\sigma\in T_{n}$, because each set of diameter at most $\alpha$ is contained in such a set $\sigma]$. So, we have $N_{\alpha}(X)=|S_{n}|$. It follows that
(1.4) $\underline{\dim}_{B}([S])$ $=$ $\liminf_{n\to\infty}\frac{\log|S_{n}|}{\log|T_{n}|}$
(1.5) $\overline{\dim}_{B}([S])$ $=$ $\limsup_{n\to\infty}\frac{\log|S_{n}|}{\log|T_{n-1}|},$
using that $\log(1/\alpha)\leq\log|T_{n}|$ in (1.4) and $\log|T_{n-1}|<\log(1/\alpha)$ in (1.5).
1.3. Box dimensions versus Hausdorff/packing dimension
By $V^{*}$ one denotes the set of tuples over a set $V$, seen as a rooted tree where the successor relation is given by appending one element of $V$. A subtree of $V^{*}$ is a nonempty subset closed under initial segments. We will always view a tree $T$ as a subtree of $\mathbb{N}$$^{*}$. Thus, $T_{k}$ consists of the strings in $T$ that have length $k$.
Definition 1.1.
We say that a tree $S\subseteq{\mathbb{N}}^{*}$ is level-wise uniformly branching if for each $n$, each $\sigma\in S_{n}$ has the same number of successors on $S$.
Example 1.2.
Let $T=\{0,1\}^{*}$. Let $S$ be a level-wise uniformly branching subtree of $T$ such that for each even $k$, for each $n\in[4^{k},4^{k+1})$, strings at level $n$ have one successor, and for each odd $k$, for each $n\in[4^{k},4^{k+1})$, strings at level $n$ have two successors. Then $\underline{\dim}_{B}([S])\leq 1/4$ and $\overline{\dim}_{B}([S])\geq 3/4$.
For background on Hausdorff dimension and also packing dimension $\dim_{P}$, see for instance the first two chapters of [4]. Here we merely review the pertinent definitions. Let $X$ be a bounded set in a metric space $M$.
For $r\geq 0$, by $H^{r}(X)$ one denotes the $r$-dimensional Hausdorff measure
(1.6) $H^{r}(X)=\lim_{\alpha\to 0^{+}}(\inf_{(E_{i})}\sum_{i\in I}\operatorname{diam}(E_{i})^{r}),$
where the infimum is taken over all covers $(E_{i})_{i\in I}$ of $X$ by sets of diameter at most $\alpha$. By definition, $\dim_{H}X$ is the supremum of the $r$ such that $H^{r}(X)>0$.
For $r\geq 0$, by $P^{r}(X)$ one denotes the $r$-dimensional packing measure
(1.7) $P^{r}(X)=\inf_{\mathcal{U}}\left\{\sum_{U\in\mathcal{U}}\lim_{\alpha\to 0^{+}}(\sup\sum_{i\in I}\operatorname{diam}(E_{i}^{U})^{r})\right\},$
where the infimum is taken over all countable coverings $\mathcal{U}$ of $X$, and the inner suprema are taken over all countable collections of disjoint open balls $(E_{i}^{U})_{i\in I}$ with center in $U$ and diameter at most $\alpha$. Now $\dim_{P}(X)$ is the supremum of the $r$ such that $P^{r}(X)>0$. It is known that
(1.8) $\dim_{H}(X)\leq\underline{\dim}_{B}(X)\text{ and }\dim_{P}(X)\leq\overline{\dim}_{B}(X)$
for any bounded subset $X$ of a metric space $M$ [4, p. 72].
2. Coincidences of dimensions
Let $S$ be a level-wise uniformly branching subtree of $T$. View $[S]$ as a subset of the metric space $[T]$. We show that the Hausdorff and lower box dimensions of $[S]$ coincide, and the packing and upper box dimensions also coincide. This section gives direct proofs of these facts. The final section provides alternative proofs using an algorithmic theory, in particular Martin-Löf randomness on $[S]$ and the point-to-set principles in metric spaces due to Lutz et al. [9]. Those proofs directly obtain the equalities from the point-to-set principles, rather than relying on the inequalities (1.8).
Theorem 2.1.
Let $S\subseteq T$ be trees such that each node on $T$ has at least two successors. Suppose that $S$ is level-wise uniformly branching. Then in the metric space $[T]$ we have
$\dim_{H}[S]$ $=$ $\underline{\dim}_{B}[S],\text{ and}$
$\dim_{P}[S]$ $=$ $\overline{\dim}_{B}[S].$
Proof.
It suffices to show the inequalities $\geq$”, as the inequalities $\leq$” always hold by (1.8). We may assume that $T$ is a subtree of $\mathbb{N}$$^{*}$.
1. For the inequality $\dim_{H}[S]\geq\underline{\dim}_{B}[S]$, we may assume $\underline{\dim}_{B}[S]>0$. It suffices to show that $H^{r}[S]\geq 1$ for each $r<\underline{\dim}_{B}[S]$ such that $r>0$.
We call a tree $X\subseteq{\mathbb{N}}^{*}$ left-closed if $\sigma\,\widehat{\ }\,i\in X$ and $k<i$ implies $\sigma\,\widehat{\ }\,k\in X$. By the definition of the ultrametric (1.1), for each tree $Y$, there is a left-closed tree $X$ such that $[Y]$ is isometric to $[X]$. So we can assume that both $T$ and $S$ are left closed.
We first provide a reduction of the type of coverings we need to consider. Write $t_{n}=|T_{n}|$ and $s_{n}=|S_{n}|$. Suppose we have a covering $(E_{i})_{i\in I}$ of $[S]$ by sets of diameter at most $\alpha$, where $t_{n-1}^{-1}>\alpha\geq t_{n}^{-1}$. In evaluating (1.6) we can assume that $\alpha=t_{n}^{-1}$. The $r$-cost of a family $(E_{i})_{i\in I}$ is by definition
$V_{r}(E_{i})_{i\in I}=\sum_{i\in I}\operatorname{diam}(E_{i})^{r}$.
In evaluating (1.6) we can further assume that each $E_{i}$ intersects $[S]$. Also, without incurring a larger $r$-cost we can replace a covering of $[S]$ by a covering with sets that contain the original sets but have no larger diameter. So we can assume that each $E_{i}$ is of the form $\sigma]$ for some $\sigma\in S$, where $|\sigma|\geq n$. Since $[S]$ is compact and each set of the form $\sigma]$ is open, we can now assume that $I$ is finite. To summarise, in evaluating (1.6), the only coverings of $[S]$ we need to consider are coverings $\{[\sigma]\colon\sigma\in F\}$ where $F\subseteq S$ is a finite prefix-free set of strings of length at least $n$. We will identify such a set $F\subseteq S$ with the covering it describes, and in particular write $V_{r}(F)$ for the $r$-cost of the covering. So $V_{r}(F)=\sum_{\rho\in F}t_{|\rho|}^{-r}$.
Claim 2.2.
Given $n$, let $F\subseteq S$ be as above. Then $V_{r}(F)\geq s_{k}t_{k}^{-r}$ for some $k\geq n$.
To verify this, we keep replacing $F$ by a covering $F^{\prime}$ of $[S]$ of no larger $r$-cost, until we reach a covering of the form $\{[\sigma]\colon\sigma\in S_{k}\}$ for some $k\geq n$, which then has the required $r$-cost $s_{k}t_{k}^{-r}$.
Let $k\geq n$ be the least length of a string in $F$. For $\sigma\in S_{k}$, write
$v(\sigma)=V_{r}\{\rho\in F\colon\,\rho\succeq\sigma\}.$
Note that $V_{r}(F)=\sum_{\sigma\in S_{k}}v(\sigma)$. Pick $\sigma\in S_{k}$ such that $v(\sigma)$ is least.
  • (a)
    If $\sigma\in F$, then $V_{r}(\{\sigma\})=t_{k}^{-r}$, whence $V_{r}(F)\geq s_{k}t_{k}^{-r}$. So we end the process.
  • (b)
    Otherwise, let $F^{\prime}=\{\eta\tau\colon\,\eta\in S_{k}\,\land\,\sigma\tau\in F\}$. Intuitively, above each $\eta\in S_{k}$, put a copy of what $F$ contains above $\sigma$. Since $S$ is level-wise uniformly branching and left-closed, we have $F^{\prime}\subseteq S$ and $F^{\prime}$ is a covering of $[S]$. Also, $V_{r}(F^{\prime})\leq V_{r}(F)$, each string in $F^{\prime}$ has a length greater than $k$, $F^{\prime}$ is prefix-free, and the maximum length of strings in $F^{\prime}$ is at most the one of $F$. So we can continue the process with $F^{\prime}$ instead of $F$.
After finitely many steps the process ends at (a), delivering the covering of $[S]$ of the required form. This shows the claim.
Since $r<\underline{\dim}_{B}[S]$, there is $K\in{\mathbb{N}}$ such that $\log s_{k}/\log t_{k}\geq r$ for each $k\geq K$, and hence $s_{k}t_{k}^{-r}\geq 1$. For $\alpha<t_{K}^{-1}$, if $(E_{i})_{i\in I}$ is a covering of $[S]$ such that $\operatorname{diam}(E_{i})\leq\alpha$ for each $i\in I$, then by the reductions and the claim we have $V_{r}(E_{i})_{i\in I}\geq s_{k}t_{k}^{-r}$ for some $k\geq K$, and hence $V_{r}(E_{i})_{i\in I}\geq 1$. Thus $H^{r}[S]\geq 1$ as required.
2. $\dim_{P}[S]\geq\overline{\dim}_{B}[S]$. Our proof is based on a result of Tricot [14] (also see [4, Lemma 2.8.1(i)]). By that result, since $[S]$ as a metric space is complete,
$\dim_{P}[S]\geq\inf\{\overline{\dim}_{B}(V)\colon V\neq\emptyset\text{ open in }[S]\};$
By the monotonicity of $\overline{\dim}_{B}$, we can, without changing the value, take the infimum over all sets $V$ of the form $\sigma]\cap[S]$ where $\sigma\in S$. Fix such a $\sigma$, and write $r=|\sigma|$. Since $S$ is level-wise uniformly branching, for $n\geq r$ the $n$-th level of the tree for $\sigma]\cap[S]$ has $s_{n}/s_{r}$ elements. So, using (1.5) we have
$\overline{\dim}_{B}(V)=\limsup_{n}\frac{\log(s_{n}/s_{r})}{\log t_{n-1}}=\limsup_{n}\frac{\log(s_{n})}{\log t_{n-1}}=\overline{\dim}_{B}[S].$
The following example shows that some hypothesis on $S$, such as being level-wise uniformly branching, is necessary in both equalities of Theorem 2.1.
Example 2.3.
Let $T=2^{<\omega}$ be the full binary tree. For a string $\sigma\in T$, let $k_{\sigma}=\min\{i\colon\,i=|\sigma|\lor\sigma_{i}=1\}$. Let
$S=\{\sigma\in T\colon\,\sigma_{r}=0\text{ for each }r\text{ such that }|\sigma|>r\geq 2k_{\sigma}\}$.
Clearly $[S]$ is countable, and hence $\dim_{H}[S]=\dim_{P}[S]=0$. In contrast, since $2\cdot 2^{n/2}\geq|S_{n}|\geq 2^{n/2}$, we have $\underline{\dim}_{B}[S]=\overline{\dim}_{B}[S]=1/2$.
3. Dimensions of closed subgroups of profinite groups
Let the topological group
[…]
be the inverse limit of an inverse system of discrete finite groups, where the maps $p_{n}\colon L_{n+1}\to L_{n}$ are epimorphisms, and $L_{0}$ is the trivial group. The group $G$ is compact and has a compatible ultrametric, where the distance between distinct elements $x,y$ of $G$ is $1/|L_{n}|$ for the largest $n$ such that the natural projections of $x$ and $y$ into $L_{n}$ are equal. Barnea and Shalev [2, Thm. 2.4] showed that the Hausdorff dimension of a closed subgroup $U$ of $G$ equals its lower box dimension. They relied on Abercrombie [1, Prop. 2.6] for the nontrivial part, that the lower box dimension is no larger than the Hausdorff dimension. Abercrombie proved this via means specific to profinite groups, such as generation of subgroups and the Haar measure. We will show in Theorem 3.1 that this is a purely geometric fact: it does not rely on the group structure, but merely exploits, via Theorem 2.1, the regularity of a description of $U$ as a subtree of the tree given by the inverse system. We also obtain the dual fact that the packing dimension of a closed subgroup equals its upper box dimension.
Define a tree $T$ such that the $n$-th level is $T_{n}=L_{n}$. A string $\eta\in T_{n+1}$ is a successor of $\sigma\in T_{n}$ if $p_{n}(\eta)=\sigma$. Each $\sigma\in T_{n}$ has $|L_{n+1}|/|L_{n}|$ successors. So $T$ is level-wise uniformly branching. There is a canonical isometry between $G$ and $[T]$: map an element $x$ of $G$ to the path $f_{x}$ such that $f_{x}(n)=r$ if the natural projection map $G\to L_{n}$ sends $x$ to $r$.
Theorem 3.1 ([2], Thm. 2.4 for the first line of equations).
[…]
$\dim_{H}(U)=\underline{\dim}_{B}(U)=\liminf_{n}\frac{\log|U_{n}|}{\log|L_{n}|}$
$\dim_{P}(U)=\overline{\dim}_{B}(U)=\limsup_{n}\frac{\log|U_{n}|}{\log|L_{n-1}|}.$
Proof.
[…]
We end the section with some discussion of fractal dimensions, in particular in profinite groups. The following example shows that $\underline{\dim}_{B}(U)<\overline{\dim}_{B}(U)$ is possible.
Example 3.2 (essentially [2], Lemma 4.1).
Let $E$ be the Cantor space $C_{2}^{\omega}$ with symmetric difference $\Delta$ as the group operation. Let the inverse system be given by $L_{n}=C_{2}^{n}$, with the canonical map $p_{n}$ that omits the first bit in a binary string of length $n+1$. For each pair of reals $\alpha,\beta$ with $0\leq\alpha\leq\beta\leq 1$ there is a closed subgroup $U$ of $G$ such that $\underline{\dim}_{B}(U)=\alpha$ and $\overline{\dim}_{B}(U)=\beta$.
Proof.
Let $R\subseteq{\mathbb{N}}$ be a set with lower [upper] density $\alpha$ $\beta$]. Let $H$ be the subgroup of sequences that have a 0 in positions outside $R$. We have $|S_{n}|=2^{|R\cap n|}$, $|T_{n}|=2^{n}$. Thus $\underline{\dim}_{B}(U)=\alpha$ and $\overline{\dim}_{B}(U)=\beta$ by (1.2) and (1.3), respectively. ∎
If $G$ is a $p$-adic analytic pro-$p$ group [5], then by [2, Thm 1.1],
$\dim_{H}(U)=\dim_{\text{manifold}}(U)/\dim_{\text{manifold}}(G)$,
where $\dim_{\text{manifold}}(X)$ is the dimension of a closed set $X$ in the $p$-adic manifold underlying $G$. Barnea and Shalev [2, end of Section 3] assert that the proof also works for the upper box dimension in place of $\dim_{H}$, so the same expression holds. Thus all the fractal dimensions agree for closed subgroups of $G$; in particular, $\dim_{H}$ and $\dim_{P}$ agree.
The spectrum of a profinite group is the set of Hausdorff dimensions of closed subgroups. The group $E$ given in Example 3.2 has the largest possible spectrum $[0,1]$. Note that $E$ is not (topologically) finitely generated. Recent research addresses questions about which spectra can occur for finitely generated groups, with particular constraints on the groups and the type of subgroups. For instance, de las Heras and Klopsch [6] proved that there is a 2-generated pro-$p$ group $G$ such that with respect to any of the five standard ways for pro-$p$ groups to represent it as an inverse limit, each real in $[0,1]$ occurs as the dimension of a normal closed subgroup.
4. Algorithmic dimension of points
The remainder of the paper is devoted to the algorithmic theory of fractal dimensions. The present section reviews the Kolmogorov complexity of points in metric spaces, and describes the point-to-set principle for both Hausdorff and packing dimension in that context. We mostly follow Lutz, Lutz and Mayordomo [9]. However, we restrict ourselves to the standard gauge family $\theta_{s}(\alpha)=\alpha^{s}$, which has a precision family in their sense [9, Observation 2.1]. For further background see [11].
Convention 4.1.
All logarithms will be taken with respect to base $2$.
Suppose one is given a metric space $M$ together with a countable dense set $D\subseteq M$ and an onto function $r\colon\{0,1\}^{*}\to D$. For $w\in D$ one lets
(4.1) $C^{M,r}(w)=\min\{C(u)\colon r(u)=w\},$
where $C(u)$ is the usual Kolmogorov complexity of a binary string $u$. (See Li and Vitanyi [7], the standard reference for Kolmogorov complexity.)
Definition 4.2 ([9], (3.1) and (3.2)).
For $x\in M$ and any $\alpha>0$, one defines the Kolmogorov complexity of a point $x$ at precision $\alpha$ by
(4.2) $C_{\alpha}(x)=\min\{C^{M,r}(w)\colon w\in D\,\land\,d(x,w)\leq\alpha\}.$
The algorithmic dimension of $x\in M$ is
(4.3) $\textit{dim}(x):=\liminf_{\alpha\to 0^{+}}C_{\alpha}(x)/\log(1/\alpha).$
The strong algorithmic dimension of $x\in M$ is
(4.4) $\textit{Dim}(x):=\limsup_{\alpha\to 0^{+}}C_{\alpha}(x)/\log(1/\alpha).$
These definitions can be readily relativised to an oracle set $A\subseteq{\mathbb{N}}$: in (4.1), replace $C(u)$ by $C^{A}(u)$, the Kolmogorov complexity of a string $u$ as measured by a universal machine equipped with $A$ as an oracle. Two point-to-set principles [9, Thms 4.1,4.2] determine the Hausdorff and packing dimension of a set $X\subseteq M$, respectively, in terms of the relativised algorithmic dimensions of the points in it:
(4.5) $\dim_{H}(X)$ $=$ $\min_{A}\sup_{x\in X}\textit{dim}^{A}(x),$
(4.6) $\dim_{P}(X)$ $=$ $\min_{A}\sup_{x\in X}\textit{Dim}^{A}(x)$
Definition 4.3.
Fix some computable injection $\pi\colon{\mathbb{N}}^{*}\to\{0,1\}^{*}$, say
$(n_{0},\ldots,n_{k-1})\mapsto 0^{n_{0}}1\ldots 0^{n_{k-1}}1$.
The plain descriptive (Kolmogorov) complexity $C(y)$ of $y\in{\mathbb{N}}^{*}$ is defined to be $C(\pi(y))$; in a similar way one defines $K(y)$, $C(y\mid n)$, etc.
The choice of $\pi$ will only affect these values up to a constant. It will be clear from the context whether $C(y)$ is taken to be evaluated for $y\in{\mathbb{N}}^{*}$ or for $y\in\{0,1\}^{*}$. There are now two interpretations of the expression $C(y)$ for a binary string $y$; however, they agree up to a constant.
Recall that throughout, $T\subseteq{\mathbb{N}}^{*}$ is a finitely branching tree such that each node has at least two successors. By $T_{n}$ we denote the set of strings in $T$ of length $n$. The space $[T]$ carries the ultrametric given by (1.1). For real-valued expressions $E,F$ depending on the same parameters, we write $E\leq^{+}F$ for $E\leq F+\mathcal{O}(1)$, and $E=^{+}F$ for $E\leq^{+}F\leq^{+}E$.
Suppose $T$ is computable. We describe a computable structure on the metric space $[T]$ in the usual sense of computable analysis [3]. That is, we describe a function $r\colon\{0,1\}^{*}\to D$ as above such that the distance of $r(u)$ and $r(v)$ is a computable real, uniformly in strings $u,v$. Let $\underline{r}\colon\{0,1\}^{*}\to T$ be the computable bijection that is an isomorphism between the length lexicographical orderings on $\{0,1\}^{*}$ and $T$. Given a binary string $u$, let $r(u)$ be a path on $[T]$ extending the string $\underline{r}(u)$. Since $T$ is computable, one can choose $r(u)$ to be uniformly computable in $u$. Clearly the distances are computable reals as required above.
We next relate $C_{\alpha}(x)$, the Kolmogorov complexity of a point at precision $\alpha$ in (4.2), to the initial segment complexity of paths on $[T]$.
Proposition 4.4.
Suppose that $T\subseteq{\mathbb{N}}^{*}$ is a computable tree such that each node has at least two successors. Let $r\colon\{0,1\}^{*}\to[T]$ define a computable structure as above. Let $f\in[T]$. For each $\alpha$ such that $|T_{n-1}|^{-1}>\alpha\geq|T_{n}|^{-1}$, we have
$C(f\!\upharpoonright_{n})-2\log n\leq^{+}C(f\!\upharpoonright_{n}\mid n)\leq^{+}C_{\alpha}(f)\leq^{+}C(f\!\upharpoonright_{n})$.
Proof.
The leftmost inequality is standard [7]. For the middle inequality, suppose $w\in D=\mathrm{range}(r)$ and $d(w,f)\leq\alpha$. Given $u\in\{0,1\}^{*}$ such that $r(u)=w$ and $n\in{\mathbb{N}}$, we can compute $w\!\upharpoonright_{n}$, which equals $f\!\upharpoonright_{n}$ since $|T_{n-1}|^{-1}>\alpha$. Thus $C(f\!\upharpoonright_{n}\mid n)\leq^{+}C(w)$. Since $w$ was arbitrary this shows $C(f\!\upharpoonright_{n}\mid n)\leq^{+}C_{\alpha}(f)$.
For the rightmost inequality, given a string $y=f\!\upharpoonright_{n}\in T$, compute the string $u\in\{0,1\}^{*}$ such that $\underline{r}(u)=y$. Since $d(f,r(u))\leq\alpha$, this shows $C_{\alpha}(f)\leq^{+}C(u)\leq^{+}C(y)$. ∎
The algorithmic dimensions $\textit{dim}(f)$ and $\textit{Dim}(f)$ (see Def. 4.2) of a path $f\in[T]$ can be expressed in terms of the descriptive complexity of its initial segments. This is well-known at least for the case of a binary tree $T$, going back to [10].
Proposition 4.5.
Let $T$ be a computable tree such that each string on $T$ has at least two successors on $T$. For any $f\in[T]$ we have
$\textit{dim}(f)=\liminf_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n}|}=\liminf_{n}\frac{K(f\!\upharpoonright_{n})}{\log|T_{n}|}$
$\textit{Dim}(f)=\limsup_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n-1}|}=\limsup_{n}\frac{K(f\!\upharpoonright_{n})}{\log|T_{n-1}|}$
Proof.
For the left equalities, recall that for $|T_{n-1}|^{-1}>\alpha\geq|T_{n}|^{-1}$ we have $C(f\!\upharpoonright_{n})-2\log n\leq^{+}C_{\alpha}(f)\leq^{+}C(f\!\upharpoonright_{n})$ by Proposition 4.4. By hypothesis $\log|T_{n}|\geq n$, and so $\lim_{n}\log n/\log|T_{n}|=0$. Therefore
$\liminf_{\alpha\to 0^{+}}C_{\alpha}(f)/\log(1/\alpha)$ $=$ $\liminf_{n\to\infty}C(f\!\upharpoonright_{n})/\log|T_{n}|$
$\limsup_{\alpha\to 0^{+}}C_{\alpha}(f)/\log(1/\alpha)$ $=$ $\limsup_{n\to\infty}C(f\!\upharpoonright_{n})/\log|T_{n-1}|$
For the right equalities, note that for each $n$ and each $y\in T_{n}$, we have $C(y)\leq^{+}K(y)\leq^{+}C(y)+K(n)\leq^{+}C(y)+2\log n$. ∎
5. Martin-Löf randomness and Levin-Schnorr Theorem on $[S]$.
Definition 5.1.
For a finitely branching tree $S$ without leaves, the path space $[S]$ carries the uniform measure $\mu_{S}$ determined by the conditions that
  • $\mu_{S}[S]=1$
  • $\mu_{S}([\sigma\,\widehat{\ }\,i]\cap[S])=\mu_{S}([\sigma]\cap[S])/k_{\sigma}$ where $k_{\sigma}$ is the number of successors of $\sigma$ on $S$.
Note that if $S$ is level-wise uniformly branching, then $\mu_{S}([\sigma])=|S_{n}|^{-1}$ for each $\sigma\in S_{n}$.
If $S$ and the function $\sigma\mapsto k_{\sigma}$ are computable, then $\mu_{S}$ is computable in the usual sense of [8]. Thus, algorithmic test notions such as Martin-Löf tests (ML-tests) can be used for $[S]$ in place of Cantor space. Also, one can generalise the Levin-Schnorr Theorem (see, e.g., [12, 3.2.9]) to $[S]$. For $X\subseteq S$ let $X]^{\prec}_{S}=\{f\in S\colon\exists n\,f\!\upharpoonright_{n}\in X\}$ be the open subset of $[S]$ described by $X$. Let
$R_{S,b}=[\{\sigma\in S\colon K(\sigma)\leq-\log\mu_{S}[\sigma]-b\}]^{\prec}_{S}.$
Our version of the Levin-Schnorr theorem is as follows.
Lemma 5.2.
$(R_{S,b})_{b\in{\mathbb{N}}}$ is a universal Martin-Löf test on $[S]$. Thus, it is a ML-test, and $f\in[S]$ is ML-random iff there is $b\in{\mathbb{N}}$ such that
$K(f\!\upharpoonright_{n})>-\log\mu_{S}[f\!\upharpoonright_{n}]-b$ for each $n\in{\mathbb{N}}$.
Proof.
Clearly the sets $R_{S,b}$ are $\Sigma^{0}_{1}$ uniformly in $b$. Given $b$, let $D$ be the set of prefix minimal strings $\sigma\in S$ such that $K(\sigma)\leq-\log\mu([\sigma])-b$. Then
$1\geq\sum_{\sigma\in D}2^{-K(\sigma)}\geq\sum_{\sigma\in D}2^{\log\mu([\sigma])+b}.$
This implies that $\mu_{S}[D]^{\prec}_{S}\leq 2^{-b}$.
Universality is shown as in the proof for Cantor space $2^{{\mathbb{N}}}$; see [12, 3.2.9]. To adapt that proof, one first needs to write down a machine existence theorem for machines with binary input that output strings in $\mathbb{N}$$^{*}$. As this is well-known in principle, we omit the details. ∎
While we are mainly interested in the case that $[S]$ is uncountable, the argument above also works for countable $[S]$. In that case, only the empty set is null, so every path is ML-random on $[S]$.
6. Algorithmic dimension versus box dimension
We show that in the case that a computable subtree $S$ of $T$ is level-wise uniformly branching, the lower/upper box dimension of $[S]$ equals the (strong) algorithmic dimension of a ML-random $f\in[S]$. To provide the desired proof of Theorem 2.1 using the point-to-set principles, this fact will be used in relativised form.
Proposition 6.1.
(i) Let $T$ be a computable tree as above, and let $S$ be a computable subtree of $T$. For each $f\in[S]$, we have
$\textit{dim}(f)$ $\leq$ $\underline{\dim}_{B}([S]),\text{ and}$
$\textit{Dim}(f)$ $\leq$ $\overline{\dim}_{B}([S]).$
(ii) Suppose $S$ is level-wise uniformly branching. Then the equalities hold in the case that $f$ is Martin-Löf random in $[S]$ with respect to $\mu_{S}$.
Proof.
Throughout we will use (1.4),(1.5) and Prop. 4.5. Let $S_{\leq n}=\bigcup_{k\leq n}S_{k}$.
(i) Let $\eta_{S}\colon S\to\{0,1\}^{*}$ be the (computable) isomorphism between the length-lexicographical orderings on $S$ and on $\{0,1\}^{*}$. Recall that to define $C(y)$ and $K(y)$ for $y\in{\mathbb{N}}^{*}$, we fixed a computable injection $\pi\colon{\mathbb{N}}^{*}\to\{0,1\}^{*}$. Also recall that logarithms are taken in base $2$.
Claim. Let $y\in S_{n}$. Then $C(y)\leq^{+}\log|S_{\leq n}|$.
To verify this, let $x=\eta_{S}(y)$. We have $|x|\leq\lceil\log|S_{\leq n}|\rceil$. Since $x$ is binary, $C(x)\leq^{+}|x|$. Since the function $\pi\circ\eta_{S}^{-1}$ is computable, we have $C(y)=C(\pi(\eta_{S}^{-1}(x)))\leq^{+}C(x)$. The claim follows.
To verify the inequalities (i), let $y=f\!\upharpoonright_{n}$. Note that
$\log|S_{\leq n}|\leq\log(n\cdot|S_{n}|)\leq^{+}\log n+\log|S_{n}|$.
So by the claim, $C(y)\leq^{+}\log n+\log|S_{n}|$. Using that $\log n=o(\log|T_{n}|)$,
$\dim(f)$ $=$ $\liminf_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n}|}$
$\leq$ $\liminf_{n}\frac{\log n+\log|S_{n}|}{\log|T_{n}|}=\underline{\dim}_{B}[S].$
Similarly,
$\operatorname{\textit{Dim}}(f)$ $=$ $\limsup_{n}\frac{C(f\!\upharpoonright_{n})}{\log|T_{n-1}|}$
$\leq$ $\limsup_{n}\frac{\log n+\log|S_{n}|}{\log|T_{n-1}|}=\overline{\dim}_{B}[S].$
(ii) If $f$ is ML random then $f$ passes the universal test $(R_{S,b})_{b\in{\mathbb{N}}}$ in Lemma 5.2. Thus, using that $S$ is level-wise uniformly branching, we have
$\log|S_{n}|=^{+}-\log\mu_{S}[f\!\upharpoonright_{n}]\leq^{+}K(f\!\upharpoonright_{n})$.
The equalities now follow from Proposition 4.5. ∎
Proof of Th. 2.1 using the point-to-set principle.
Suppose that $E\subseteq{\mathbb{N}}$ is an oracle such that $S\leq_{T}E$, $T\leq_{T}E$, and $E$ can compute from $n$ the number of successors for $S$ at level $n$. Note that the uniform measure $\mu_{S}$ on $[S]$ is $E$-computable.
We show $\dim_{H}[S]=\underline{\dim}_{B}[S]$. By the point-to-set principle for Hausdorff dimension (4.5), it suffices to show that $\min_{A}\sup_{f\in[S]}\textit{dim}^{A}(f)=\underline{\dim}_{B}([S])$. We have the inequality $\leq$” by Prop. 6.1(i) relativised to $E$. For the converse inequality, note that in taking the minimum on the left hand side, we may assume that the oracle $A$ computes $E$, because replacing $A$ by $A\oplus E$ does not increase $\textit{dim}^{A}(f)$. For each such $A$, if we choose $f$ to be MLA-random in $[S]$, by Prop. 6.1(ii) we have $\textit{dim}^{A}(f)=\underline{\dim}_{B}(S)$.
The second equality in Th. 2.1 is obtained in a similar fashion, using the point-to-set principle for packing dimension (4.6) instead. ∎